# An Open and Large-Scale Dataset for Multi-Modal Climate Change-aware Crop Yield Predictions

### v1

@@ 1. Introduction
Precise crop yield prediction is essential for early agricultural planning (Khaki et al., 2021), timely management policy adjustment (Turchetta et al., 2022), informed financial decision making (Ansarifar et al., 2021), and national food security (Mourtzinis et al., 2021). Recent advancements in deep neural networks (DNNs) have achieved impressive performance across various domains (Krizhevsky et al., 2012; He et al., 2016; Vaswani et al., 2017; Dosovitskiy et al., 2021; He et al., 2021; Chen et al., 2022c; Lin et al., 2022; Mo et al., 2022; Chen et al., 2022b; Ma et al., 2022; Feng et al., 2023; Lin et al., 2023b; Ruan et al., 2023; Li et al., 2024; Mo et al., 2024b; Zhang et al., 2024; Song et al., 2024). Building upon these advancements, plenty of studies have employed spatial-temporal DNNs (Wang et al., 2022; Ma et al., 2023; Li et al., 2023b; Liu et al., 2023; Wang et al., 2023b; Li et al., 2022b; Ma et al., 2024; Mo et al., 2024a; Li et al., 2022a; He et al., 2022b) to predict crop yields with increased timeliness and precision (Khaki et al., 2020, 2021; Garnot and Landrieu, 2021; Wu et al., 2021; Cheng et al., 2022; Fan et al., 2022; Lin et al., 2023a). However, they often applied their personally curated and limit-sized datasets, with somewhat mediocre prediction performance. There is an urgent need for new large-scale and deep learning-ready datasets tailored specifically for wide use in crop yield predictions.
@@ 4.4. Improving the Generalization Capabilities of DNNs
Self-supervised learning (SSL) techniques (Chen et al., 2020; Bao et al., 2022; He et al., 2022a; Zhou et al., 2023; Zhang et al., 2023; Wang et al., 2023a) have significantly advanced the generalization capabilities of deep neural networks (DNNs), especially in vision transformers (ViTs). Our CropNet dataset with a total size of over $2$ TB of data can benefit both deep-learning and agricultural communities by providing large-scale visual satellite imagery and numerical meteorological data for pre-training DNNs. To exhibit the applications of our CropNet dataset to self-supervised pre-training, we adopt the MMST-ViT for crop yield predictions by considering three scenarios, i.e., MMST-ViT without the SSL technique (denoted as “w/o SSL”), MMST-ViT with the SSL technique in MAE (denoted as “MAE”), and MMST-ViT with the multi-modal SSL technique proposed in (Lin et al., 2023a) (denoted as “MM-SSL”). Figure 8 illustrates the performance results for four crop types under three performance metrics of interest (i.e., RMSE, R2, and Corr). We discover that without the SSL technique (i.e., the gray line), the MMST-ViT model exhibits limitations in generalization capabilities, resulting in suboptimal crop yield prediction performance across all tested scenarios. Besides, pre-training MMST-ViT with the SSL technique in MAE (i.e., the blue line) improves its performance results (compared to the “w/o SSL”), with decreased RMSE values by $3.8$, $9.6$, $1.3$, and $1.7$ for corn, cotton, soybeans, and winter wheat predictions, respectively. This statistical evidence confirms that our CropNet dataset can improve the generalization capabilities in vision models. Furthermore, MMST-ViT with the multi-modal SSL technique (i.e., the green line) achieves the best performance results under all scenarios. In comparison to the “w/o SSL” scenario, it decreases RMSE values by $6.4$, $18.3$, $2.6$, and $3.6$, respectively, for predicting corn, cotton, soybeans, and winter wheat. The effectiveness of the multi-modal SSL technique may stem from its ability to integrate visual satellite imagery with numerical meteorological data found in the CropNet dataset. This integration enhances the generalization capabilities of the MMST-ViT model by improving its ability to effectively discern the influence of weather conditions on crop growth patterns during the pre-training phase.

### v2

@@ 1. Introduction
Precise crop yield prediction is essential for early agricultural planning (Khaki et al., 2021), timely management policy adjustment (Turchetta et al., 2022), informed financial decision making (Ansarifar et al., 2021), and national food security (Mourtzinis et al., 2021). Recent advancements in deep neural networks (DNNs) have achieved impressive performance across various domains (Krizhevsky et al., 2012; He et al., 2016; Vaswani et al., 2017; Dosovitskiy et al., 2021; He et al., 2021; Wang et al., 2022a; Chen et al., 2022c; Lin et al., 2022; Mo et al., 2022; Chen et al., 2022b; Ma et al., 2022; Feng et al., 2023; Lin et al., 2023b; Ruan et al., 2023; Li et al., 2024; Mo et al., 2024b; Zhang et al., 2024; Zhan et al., 2024; Song et al., 2024). Building upon these advancements, plenty of studies have employed spatial-temporal DNNs (Zhan et al., 2019; Wang et al., 2022b; Ma et al., 2023; Li et al., 2023b; Liu et al., 2023; Wang et al., 2023c; Li et al., 2022b; Ma et al., 2024; Mo et al., 2024a; Li et al., 2022a; He et al., 2022b) to predict crop yields with increased timeliness and precision (Khaki et al., 2020, 2021; Garnot and Landrieu, 2021; Wu et al., 2021; Cheng et al., 2022; Fan et al., 2022; Lin et al., 2023a). However, they often applied their personally curated and limit-sized datasets, with somewhat mediocre prediction performance. There is an urgent need for new large-scale and deep learning-ready datasets tailored specifically for wide use in crop yield predictions.
@@ 4.4. Improving the Generalization Capabilities of DNNs
Self-supervised learning (SSL) techniques (Chen et al., 2020; Bao et al., 2022; He et al., 2022a; Zhou et al., 2023; Zhang et al., 2023; Wang et al., 2023b, a) have significantly advanced the generalization capabilities of deep neural networks (DNNs), especially in vision transformers (ViTs). Our CropNet dataset with a total size of over $2$ TB of data can benefit both deep-learning and agricultural communities by providing large-scale visual satellite imagery and numerical meteorological data for pre-training DNNs. To exhibit the applications of our CropNet dataset to self-supervised pre-training, we adopt the MMST-ViT for crop yield predictions by considering three scenarios, i.e., MMST-ViT without the SSL technique (denoted as “w/o SSL”), MMST-ViT with the SSL technique in MAE (denoted as “MAE”), and MMST-ViT with the multi-modal SSL technique proposed in (Lin et al., 2023a) (denoted as “MM-SSL”). Figure 8 illustrates the performance results for four crop types under three performance metrics of interest (i.e., RMSE, R2, and Corr). We discover that without the SSL technique (i.e., the gray line), the MMST-ViT model exhibits limitations in generalization capabilities, resulting in suboptimal crop yield prediction performance across all tested scenarios. Besides, pre-training MMST-ViT with the SSL technique in MAE (i.e., the blue line) improves its performance results (compared to the “w/o SSL”), with decreased RMSE values by $3.8$, $9.6$, $1.3$, and $1.7$ for corn, cotton, soybeans, and winter wheat predictions, respectively. This statistical evidence confirms that our CropNet dataset can improve the generalization capabilities in vision models. Furthermore, MMST-ViT with the multi-modal SSL technique (i.e., the green line) achieves the best performance results under all scenarios. In comparison to the “w/o SSL” scenario, it decreases RMSE values by $6.4$, $18.3$, $2.6$, and $3.6$, respectively, for predicting corn, cotton, soybeans, and winter wheat. The effectiveness of the multi-modal SSL technique may stem from its ability to integrate visual satellite imagery with numerical meteorological data found in the CropNet dataset. This integration enhances the generalization capabilities of the MMST-ViT model by improving its ability to effectively discern the influence of weather conditions on crop growth patterns during the pre-training phase.